3X+11=9X SOLVE FOR X
Answer by checkley77(12844)

3X+11=9X
3X-9X=-11
-6X=-11
X=-11/-6
X=11/6 ANS.
PROOF:
3*11/6+11=9*11/6
33/6+11=99/6
5.5+11=16.5
16.5=16.5